The following information relates to Hogs Back Falls Ltd.’s inventory transactions during the month of February.

Units Cost/Unit Amount
Feb. 1 Beginning inventory 7,500 $23.00 $172,500
10 Sale 6,500
14 Purchase 4,000 $24.00 $96,000
18 Purchase 500 $28.00 $14,000
25 Sale 4,300
26 Purchase 4,200 $33.00 $138,600
28 Sale 3,200


All of the units sold were priced at $92.00 per unit.

Hogs Back Falls Ltd. uses the perpetual inventory system. Calculate Hogs Back’s cost of goods sold, gross margin, and ending inventory for the month of February using FIFO.

Cost of Goods Sold $
Gross Margin $
Ending Inventory $

Hogs Back Falls Ltd. uses the perpetual inventory system. Calculate Hogs Back’s cost of goods sold, gross margin, and ending inventory for the month of February using weighted-average. (Round calculations for cost per unit to 2 decimal places, e.g. 10.52 and final answers to 0 decimal places, e.g. 61,052.)  

Cost of Goods Sold $
Gross Margin $
Ending Inventory $

Which of the cost formulas would produce the higher gross margin?

The                                                                       FIFOweighted average method results in the higher gross margin.
Answer #1
Perpetual system and FIFO method :
Date Purchase Sales (at cost) Balance
Units Unit cost Total cost Units Unit cost Total cost Units Unit cost Total cost
Feb. 1 7500 23 172500
Feb. 10 6500 23 149500 1000 23 23000
Feb. 14 4000 24 96000 1000 23 23000
4000 24 96000
Feb. 18 500 28 14000 1000 23 23000
4000 24 96000
500 28 14000
Feb. 25 1000 23 23000 700 24 16800
3300 24 79200 500 28 14000
Feb. 26 4200 33 138600 700 24 16800
500 28 14000
4200 33 138600
Feb. 28 700 24 16800 2200 33 72600
500 28 14000
2000 33 66000
Total units sold = 6500 + 4300 + 3200 14000
Sales = ( Total units sold * Selling price ) = ( 14000 * 92.00 ) 1288000
Cost of goods sold = 149500 + 23000 + 79200 + 16800 + 14000 282500
Gross margin = Sales - Cost of goods sold = 1288000 - 282500 1005500
Ending inventory 72600
Perpetual system and weighted average method :
Date Purchase Sales (at cost) Balance
Units Unit cost Total cost Units Unit cost Total cost Units Unit cost Total cost
Feb. 1 7500 23 172500
Feb. 10 6500 23 149500 1000 23 23000
Feb. 14 4000 24 96000 5000 23.8 119000
Feb. 18 500 28 14000 5500 24.18 133000
Feb. 25 4300 24.18 103974 1200 24.19 29026
Feb. 26 4200 33 138600 5400 31.04 167626
Feb. 28 3200 31.04 99328 2200 31.04 68298
Cost of goods sold = 149500 + 103974 + 99328 352802
Gross margin = Sales - Cost of goods sold = 1288000 - 352802 935198
Ending inventory 68298
The FIFO method results in the higher gross margin